OpenStudy (anonymous):

Find the equation of a line perpendicular to the given line: y = 2/3 x + 8.

OpenStudy (anonymous):

two lines are perpendicular if they the product of their slopes is -1

OpenStudy (anonymous):

\[m1 * m2 = -1\]

OpenStudy (anonymous):

you are given m1 in your first equation

OpenStudy (anonymous):

\[m1 = 2/3\]

OpenStudy (anonymous):

you can divine \[m2\] from the relation \[m1*m2=-1\]

OpenStudy (anonymous):

\[y=m2*x + b\] is the equation of the perpendicular line

OpenStudy (anonymous):

in that they are perpendicular means that the lines intersect at a point